OpenStudy (anonymous):

find the y-intercepct for 2x + y=6

OpenStudy (anonymous):

When x = 0, y = 6. Y-intercept is thus (0,6) or simply 6.

OpenStudy (gw2011):

To find the y-intercept use the following equation: y=mx+b where m is the slope and b is the y-intercept Put the equation 2x+y=6 in the form of y=mx+b and you will get: y=-2x+6 Therefore, the y-intercept is 6 which corresponds to b in the equation y=mx+b